During the year, the Senbet Discount Tire Company had gross sales of $1.2 million. The company’s cost of goods sold and selling expenses were $589,000 and $242,000, respectively. The company also had notes payable of $810,000. These notes carried an interest rate of 5 percent. Depreciation was $119,000. The tax rate was 24 percent.

a.

What was the company’s net income? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, rounded to the nearest whole dollar amount, e.g., 1,234,567.)

b.

What was the company’s operating cash flow? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, rounded to the nearest whole dollar amount, e.g., 1,234,567.)
Answer #1

a) Net income = (Sales - COGS - Selling Expenses -Depreciation - Interest)*(1-Tax rate ) = (1,200,000 - 589,000 - 242,000 -11900 - 5%*810000)*(1-24%) = 159,220
Tax = (Sales - COGS - Selling Expenses -Depreciation - Interest)*Tax rate )= (1,200,000 - 589,000 - 242,000 -11900 - 5%*810000)*24% = 50,280

b) Operating Cash Flow = EBIT - Taxes + Depreciation = Sales - COGS - Selling Expenses -Taxes = 1,200,000 - 589,000 - 242,000 - 50,280 = 318,720

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well